OpenStudy (anonymous):

MEDAL + FAN!! A line has a slope of 7 and contains the point (5, −9).Which equations represent the line? Choose all answers that are correct.

OpenStudy (anonymous):

any choices?

OpenStudy (anonymous):

A. −7x + y = −44 B. 7x – y = 44 C. (y − 9) = 7(x + 5) D. (y + 9) = 7(x − 5)

OpenStudy (anonymous):

ok just a second

OpenStudy (anonymous):

y=7x-44

OpenStudy (anonymous):

so just convert give me a minute

OpenStudy (anonymous):

Thank you for the help

OpenStudy (anonymous):

wait a minute there can be many answers

OpenStudy (anonymous):

well A, C

OpenStudy (anonymous):

looks like you got this

OpenStudy (anonymous):

A, D